Problem 6-31 Equivalent annual costs The presidents executive jet is not fully utilized. You judge that its use by other officers would increase direct operating costs by only $22,000 a year and would save $100,000 a year in airline bills. On the other hand, you believe that with the increased use the company will need to replace the jet at the end of three years rather than four. A new jet costs $1.12 million and (at its current low rate of use) has a life of five years. Assume that the company does not pay taxes. All cash flows are forecasted in real terms. The real opportunity cost of capital is 10% a. Calculate the equivalent annual cost of a new jet. (Do not round intermediate calculations. Enter your answer as a negative value rounded to 2 decimal places.) Equivalent annual cost 302,702 b. Calculate the present value of the additional cost of replacing the jet one year earlier than under its current usage.(Do not round intermediate calculations. Enter your answer as a negative value rounded to 2 decimal places.) Present value c. Calculate the present value of the savings. (Do not round intermediate calculations. Enter your answer as a positive value rounded to 2 decimal places.) Present value 193,974. d. Should you try to persuade the president to allow other officers to use the plane? O No O YesPlease, I won't get a and b right. Thank you!
Answer #1
a) EAC of the new jet = 1120000*0.10*1.1^5/(1.1^5-1) = $      2,95,453.18
b) PV of the additional cost = 22000/1.1 = $          20,000.00
c) PV of the savings = 100000/1.1 = $          90,909.09
d) No